# OpticalRS-4M: Scaling Efficient Masked Autoencoder Learning on Large Remote Sensing Dataset

### v3

@@ 4.2.1 Fine-tuning Performance on Downstream Tasks
We further evaluated SelectiveMAE with the OpticalRS-4M dataset across three key downstream tasks: scene classification, object detection, and semantic segmentation. In addition to benchmarking against MAE-based approaches like SatMAE[30], ScaleMAE[33], SSL4EO[54], RVSA [18], RingMo[37], and GFM[25], we compared it with other leading visual foundation models. These encompass methods leveraging contrastive learning and enhanced backbones, such as GASSL [24], SeCo [22], TOV [28], and CACo [21], and advanced supervised learning-based approaches [56]. Notably, none of these comparative methods were specifically tailored for accelerated pre-training, and although their speed performance wasn’t reported, we anticipate they may lag behind the computationally efficient baseline (MAE). More details can be found in the Appendix.
@@
Scene Classification. We begin by evaluating the pre-trained model’s performance on the scene classification task, which offers insight into its overall representation capability without requiring additional decoders. We leverage two scene classification datasets: AID [85] and RESISC-45 [86]. Training details, including the training and test split ratio, adhere to [18, 33], with further specifics available in our code. Evaluation is based on overall accuracy (OA). Table 6 presents the results, demonstrating SelectiveMAE’s competitive performance compared to other pre-training methods on both datasets. Notably, our method enjoys a significantly faster training speed, e.g., over 2$\times$ faster than the MAE baseline. Besides, it outperforms RVSA and others when scaled to ViT-L. This demonstrates that, with our SelectiveMAE training method on the OpticalRS-4M dataset, the model can learn strong feature representations and scale efficiently.
Table 6: Results for scene classification, object detection, and semantic segmentation. “TR" represents the ratio of training data to the entire dataset.
Model Publication Backbone Sence Object Semantic
Classification Detection Segmentation
AID RESISC-45 DIOR DIOR-R LoveDA SpaceNetv1
(TR=20%/50%) (TR=10%/20%) mAP50 mAP50 mIoU mF1
SeCo [22] ICCV’21 ResNet-50 93.47/95.99 89.64/92.91 - - 43.63 77.09
GASSL [24] ICCV’21 ResNet-50 93.55/95.92 90.86/93.06 67.40 65.65 48.76 78.51
TOV [28] JSTARS’23 ResNet-50 95.16/97.09 90.97/93.79 70.16 66.33 49.70 -
CACo [21] CVPR’23 ResNet-50 90.88/95.05 88.28/91.94 66.91 64.10 48.89 77.94
SatMAE [30] NIPS’22 ViT-L 95.02/96.94 91.72/94.10 70.89 65.66 - 78.07
ScaleMAE [33] ICCV’23 ViT-L 96.44/97.58 92.63/95.04 73.81 66.47 - -
SSL4EO [54] GRSM’23 ViT-S 91.06/94.74 87.60/91.27 64.82 61.23 - -
RingMo [37] TGRS’22 Swin-B 96.90/98.34 94.25/95.67 75.90 - - -
SatLas [56] ICCV’23 Swin-B 94.96/97.38 92.16/94.70 74.10 67.59 - -
GFM [25] ICCV’23 Swin-B 95.47/97.09 92.73/94.64 72.84 67.67 - -
RVSA [18] TGRS’23 ViT-B+RVSA 97.03/98.50 93.93/95.69 75.80 68.06 51.95 -
SelectiveMAE - ViT-B 96.78/98.12 93.35/94.58 75.70 67.78 53.05 79.50
SelectiveMAE - ViT-L 97.25/98.48 94.57/95.77 77.80 70.31 54.31 79.46
@@
Horizontal & Oriented Object Detection. We utilized the well-established DIOR dataset for horizontal object detection [87] and its improved variant DIOR-R for oriented object detection [88], both comprising RGB images. Following the methodologies of previous works [18, 37], we maintained consistent experimental setups, employing Faster-RCNN [89] and Oriented-RCNN [90] as detectors for each dataset. Detailed experimental configurations are available in the accompanying code. The results are summarized in Table 6. Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [37]. When using the larger ViT-L backbone, SelectiveMAE shows enhanced performance across both detection datasets, underscoring the excellent scalability of our method.
@@
Semantic Segmentation. We further evaluate the performance of the pre-trained model on pixel-level perception tasks, particularly semantic segmentation, using two well-known remote sensing datasets: LoveDA [51] and SpaceNetv1 [52]. Our implementation closely follows [18], utilizing UperNet [91] as the segmentation framework, with mean Intersection over Union (mIOU) as the evaluation metric. Table 6 demonstrates the clear superiority of SelectiveMAE over its competitors in pixel-level semantic segmentation. The selection and reconstruction strategies employed by SelectiveMAE focus on semantically rich patches, such as the boundaries between foreground objects and background stuff, leading to better representation learning for segmentation.
@@ 5 Conclusion
In this paper, we introduce a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. Built upon it, we first curated OpticalRS-4M, a large-scale optical remote sensing dataset for unsupervised learning. Unlike previous RS datasets, OpticalRS-4M offers a larger and more diverse image set with fine-grained details relevant to downstream tasks. Benchmarking representative MIM methods on OpticalRS-4M highlights its advantages in these tasks. Then, we present SelectiveMAE to reduce the computational overhead of MIM training on large-scale RS datasets. This efficient MIM method dynamically encodes and reconstructs tokens based on their semantic richness. SelectiveMAE significantly accelerates training, demonstrating that only 40% RS image patches are needed to train a comparable MIM model. Extensive experiments demonstrate that OpticalRS-4M significantly improves classification, detection, and segmentation performance, while SelectiveMAE achieves over 2×2$\times$2 × speedup, highlighting the effectiveness and scalability of our pipeline in developing RS foundational models.

### v4

@@ 4.2 Main Results
We compare SelectiveMAE to state-of-the-art RSFMs. In addition to benchmarking against MIM-based approaches such as SatMAE [17] and ScaleMAE [56], we also compare with contrastive learning-based methods, including GASSL [1] and SeCo [49], as well as advanced supervised learning models like SatLas [5]. Notably, none of these comparative methods were specifically designed for accelerated pre-training. Although their speed performance was not reported, we expect them to lag behind the baseline MAE.
@@
Scene Classification. We first assess the pre-trained model’s performance on the scene classification task to evaluate its overall representational capability without additional decoders. We use two scene classification datasets: AID [86] and RESISC-45 [13], following training details and train-test splits as in [70, 56]. Table 5 shows that SelectiveMAE performs competitively against other pre-training methods on both datasets. Notably, our method enjoys a significantly faster training speed, e.g., over 2$\times$ faster than the MAE baseline. Additionally, it surpasses LSKNet and others when scaled to ViT-L, indicating that, with SelectiveMAE on the OpticalRS-13M dataset, the model effectively learns strong feature representations and scales efficiently.
@@
Horizontal & Oriented Object Detection. We utilized the well-established DIOR dataset for horizontal object detection [41] and its improved variant DIOR-R for oriented object detection [15]. Following the methodologies of previous works [70, 63], we maintained consistent experimental setups, employing Faster-RCNN [57] and Oriented-RCNN [88] as detectors for each dataset. The results are summarized in Table 5. Our approach, utilizing a ViT-B backbone, demonstrates competitive or superior performance compared to other methods with a Swin-B backbone, such as RingMo [63]. When using the larger ViT-L backbone, SelectiveMAE shows enhanced performance across both detection datasets, even outperforming OREOLE, which has close to 1B parameters, underscoring the excellent scalability of our method.
@@
Semantic Segmentation. We further evaluate the performance of the pre-trained model on pixel-level perception tasks, particularly semantic segmentation, using two well-known remote sensing datasets: LoveDA [76] and SpaceNetv1 [68]. Our implementation follows [70], utilizing UperNet [87] as the segmentation framework. Table 5 demonstrates the clear superiority of SelectiveMAE over its competitors in semantic segmentation tasks. SelectiveMAE focus on semantically rich patches, such as the boundaries between foreground objects and background stuff, leading to better representation learning for segmentation.
@@ 5 Conclusion
In this paper, we introduce a new pre-training pipeline for RS models, featuring the creation of a large-scale RS dataset and an efficient MIM approach. Built upon it, we first curated OpticalRS-13M, a large-scale optical remote sensing dataset for unsupervised learning. Unlike previous RS datasets, OpticalRS-13M offers a larger and more diverse image set with fine-grained details relevant to downstream tasks. Benchmarking representative MIM methods on OpticalRS-13M highlights its advantages in these tasks. Then, we present SelectiveMAE to reduce the computational overhead of MIM training on large-scale RS datasets. This efficient MIM method dynamically encodes and reconstructs tokens based on their semantic richness. SelectiveMAE significantly accelerates training, demonstrating that using only 40% RS image patches is sufficient for training a comparable MIM model. Extensive experiments show that OpticalRS-13M significantly contributes to improving classification, detection, and segmentation performance, while SelectiveMAE achieves over a 2×2$\times$2 × speedup along with GPU memory savings, highlighting the effectiveness and scalability of our pipeline in developing RS foundational models.